P71-002
P71-002: The Smaller NOT Astrometric Quasar Survey (SNAQS)
P.I.: J. P. U. Fynbo
This is a Monitoring program using ALFOSC.
Instrument setup
- ALFOSC
- Spectroscopy with 1.0/1.3 arcsec slit and Grism#4 and blocking filter #94
Observing instructions
- On the Queue page, make sure the OB-generated observing script is written to disk, or else 'compile'
- Go to directory: cd ~/obs/scripts/71-002/
and do ls -lrt
- Start the requested OB generated script
Program Tracking |
Proposal |
71-002 |
Semester |
P71: 20250401 - 20251001 |
Principal Investigator |
J. P. U. Fynbo |
Programme |
The Smaller NOT Astrometric Quasar Survey (SNAQS) |
Instrument |
ALFOSC |
Time Allocation |
|
Total Nights |
19h |
Type |
Monitoring |
Notes |
|
Total Time Spent |
08:18 (hh:mm)
|
Proposal |
Date |
Time |
Type |
Targets |
Files |
Comments |
Status |
|
71-002 |
2025-04-09 |
00:00 |
Monitoring |
|
|
100% humidity throughout the night. Could not open. |
Failed |
|
71-002 |
2025-04-18 |
01:30 |
Monitoring |
sdssJ130548+244845, sdssJ130550+230841, sdssJ133918+244619 |
ALId180156 - ALId180164, ALId180165 - ALId180173, ALId180181 - ALId180189, ALId180148 - ALId180150, ALId180151 - ALId180155 |
clear skies, the seeing was around 1.0 and 1.2"
We aborted (ctrl+c)a couple of times due to connection issues. The files from ALId180148 to ALId180155 are not usable.
For sdssJ133918+244619, despite the target being positioned clearly on the slit, the quickspec looked strange. We are setting the status as completed but let us know if the data is usable. |
Observed |
|
71-002 |
2025-04-26 |
02:29 |
Monitoring |
sdssJ124216+243332, sdssJ124721+235733, sdssJ135559+240315, sdssJ132740+332419 |
ALId260123-ALId260139,ALId260146 -ALId260163 |
Dark, clear sky and good seeing. |
Observed |
|
71-002 |
2025-04-27 |
03:08 |
Monitoring |
sdssJ125946+234908, sdssJ125941+240616, sdssJ130108+240411, sdssJ125950+235634, sdssJ134407+240709 |
ALId270093 - ALId270136 |
Clear skies and good seeing |
Observed |
|
71-002 |
2025-04-28 |
01:11 |
Monitoring |
sdssJ125549+224516, sdssJ135525+264100 |
ALId280113 - ALId280121, ALId280126 - ALId280134 |
Clear skies, and the seeing was oscillating between 1.0" and 1.1". |
Observed |
|
71-002 |
2025-04-29 |
00:00 |
Monitoring |
- |
- |
The target wasn't observed due to high humidity. The telescope was closed for the night due to
freezing conditions. |
Failed |
|
The above information is collected from the ToO/Mon reports that have been created by the observer.
Should any of the observations have been carried out using the OB system, you will find additional information
HERE
|